# Automated Tabletop Exfoliation and Identification of Monolayer Graphene Flakes

### v1

@@ Abstract
The discovery of graphene, one of the most-studied materials in condensed matter physics due to its singular mechanical, optical, and electronic properties, was enabled by manual “Scotch Tape” exfoliation. Nearly two decades later, this method is still widely used to obtain chemically-pristine flakes of graphene and other 2D van der Waals materials. Unfortunately, the yield of large, pristine flakes with uniform thickness is inconsistent. Thus, significant time and effort are required to exfoliate and locate flakes suitable for fabricating multilayer van der Waals heterostructures. Here, we describe a relatively affordable tabletop device (the ”eXfoliator”) that can reproducibly control key parameters and largely automate the exfoliation process. In a typical exfoliation run, the eXfoliator produces 3 or more large (m) high-quality monolayers, allowing new users to produce large pristine graphene monolayers at a rate comparable to manual exfoliation by an experienced user. We use an automated mapping system and computer vision algorithm to locate candidate flakes.
@@ I Introduction
Monolayer graphene, a single layer of carbon atoms arranged in a uniform honeycomb lattice with an atomic spacing of 1.42Å, has remarkable mechanical, optical, and electronic properties. Graphene has recently emerged as an exciting basis for studying strongly-correlated electron physics. Naturally occurring stacking orders of multi-layer graphene, such as Bernal bilayer graphene, ABC-trilayer graphene, and pentalayer rhombohedral graphene, can display superconductivity [1, 2, 3] and orbital magnetism[4, 5]. In addition to these configurations exfoliated from natural graphite, monolayer graphene can be artificially stacked with interlayer twists to form moiré materials with a large structural design space, including configurations with flat electronic bands and non-trivial band topology. Superconductivity and orbital magnetism[6, 7], and hints of fractional Chern insulators[8, 9], have been discovered in various graphene-based moiré materials. Motivated by the rich set of target structures based on this single material, we base our discussion and experimental demonstration on graphene, but we anticipate our methods and workflow will apply more broadly to the burgeoning family of exfoliatable van der Waals layered materials.
@@
In the Scotch tape exfoliation method introduced by Geim and Novoselov in 2004[10], natural or man-made graphite crystals –- which may consist of – graphene layers arranged in large crystallographic domains –- are repeatedly adhered to adhesive tape and cleaved apart to populate large areas of the tape with increasingly thin graphite crystals. The graphite-populated surface of the tape is then applied to a Si wafer coated with a thin SiO layer. The tape-graphite-wafer sandwich is then annealed, typically to approximately 100C, before the tape is peeled away. Graphene flakes with a distribution of thicknesses ranging down to just one layer cleave from the graphite and are left on the wafer, when adhesion to the SiO overcomes interlayer binding in the graphite.
@@
Other techniques for producing large-area monolayers from bulk graphite, including electrostatic exfoliation, gold-tape exfoliation, intercalation, and sonication continue to be developed, and graphene can also be grown over large areas by CVD. Each of these techniques has its own advantages, as discussed below, but Scotch tape exfoliation remains pre-eminent when seeking low disorder graphene-based structures in a research – as opposed to industrial production – environment. However, with the exception of CVD growth, these methods are inherently limited by the size of domains in the source graphite crystal.
@@
Electrostatic exfoliation uses high potentials between two metallic electrodes to directly cleave graphite into mono- and few-layer graphene flakes, but often results in smaller flakes with higher numbers of defects, especially in monolayers[11]. The limited size and low quality of resulting monolayers render this technique ill-suited to produce flakes destined for twisted bilayer graphene or other multilayers with delicate electronic structure.
@@
Both metal-assisted and gold-tape exfoliation techniques can produce large monolayers of TMD materials; furthermore, nickel-assisted exfoliation has been used to successfully extract single-crystal graphene monolayers from bilayer graphene grown on SiC(0001)[12]. These techniques depend on strongly adhering the target monolayer to a metal film. The monolayer is subsequently transferred to a new substrate and the metal film chemically etched, but complete removal of the metal and etchant typically cannot be accomplished without damaging the graphene. Sonication and intercalation to extract monolayer graphene leave adsorbates on the graphene and often result in flakes smaller than 1 micron[12, 13], again undesirable when planning to stack it into multilayers and study exotic electronic states.
@@
Large-area growth techniques have produced promisingly large single-crystal graphene monolayers on a variety of substrates[14, 15, 16]. Transfer of these monolayers to silicon has typically involved solvents and sometimes etchants. Dry transfer, most suitable for seeking exotic electronic states, has been rare for these grown films, despite notable successes in oxidizing a copper growth substrate then picking up graphene using exfoliated hBN[17].
@@
The design of a tabletop machine to perform Scotch tape exfoliation was primarily inspired by the observation that the Scotch tape exfoliation process takes considerable time and has inconsistent yields. We hypothesized that if the key parameters of the exfoliation process could be identified, optimal values could be determined through experiments using a purpose-built tool, and then reproduced in successive exfoliations with that tool. By translating exfoliation to a semi-automated process run on a machine, a single standardized exfoliation could produce more large single-crystal freestanding monolayers than a series of manual exfoliations, in a comparable amount of time. Furthermore, the search for optimal parameters need only be carried out once, so a sufficiently accessible design would allow individual operators to exfoliate near-optimally after a much shorter learning period than that required for manual exfoliation. Additionally, if optimal parameters existed outside the range that a human researcher could implement consistently in manual exfoliation (e.g., peeling at 0.1 mm/s or maintaining the chip at 80C during peeling), a machine might be able to achieve better results than even experienced practitioners of manual exfoliation. Finally, such a machine would free the operator to pursue other work while running an otherwise tedious or time-consuming exfoliation recipe.
In fact, based on similar considerations, others have already implemented automated exfoliation systems.
@@
A prominent example of automated exfoliation is Brookhaven’s Quantum Material Press (QPress)[18], a sophisticated cluster of almost fully automated tools capable of exfoliating, cataloguing, stacking, and storing 2D materials in an inert atmosphere. QPress has successfully automated most of the tape exfoliation process, is accessible through user proposals, and may eventually provide flakes to many collaborators. In the opposite limit, a simple automated exfoliation tool can consist of a single linear motor in tandem with a hot plate[19].
@@
The tool presented in this paper (the ”eXfoliator”) was designed to fill a middle ground, as a user-friendly machine capable of controlling a range of likely-relevant parameters during exfoliation, balancing practicality and utility as an exfoliation tool with simplicity of design and construction. The parameters of Scotch tape exfoliation targeted by the design are peel angle, peel rate, application pressure, temperature of the substrate at all stages, and dwell time under pressure or at temperature.
@@
Estimates of graphite’s interlayer van der Waals binding energy vary, but most calculated values for AB, AAA, and ABC-stacked graphite range from 40-60 meV/atom[20]. For comparison, the thermal energy of graphite at 300 (400) K is approximately 1300 (2400) J/mol, corresponding to 13.5 (25) meV/atom[21]. Because the thermal energy at accessible temperatures – as limited by the melting point of the tape’s backing – does not cross or approach the binding energy threshold, tuning the annealing or peeling temperature is assumed to primarily affect the viscoelastic behavior of the polymer tape adhesive.
@@
From the time-temperature superposition principle of polymer flow[22], adjusting the annealing temperature, application pressure, or dwell time is believed to allow the viscoelastic polymer tape adhesive to flow more freely over the graphite flakes and allow better contact before being cooled and peeled. Likewise, adjusting peel speed is assumed to tune the viscoelastic behavior of the tape adhesive instead of significantly affecting the fundamental interlayer interactions of the graphene layers. Adjusting the peel angle is believed to both change the radius of curvature of the tape and change the ratio of shearing to normal forces at the interlayer boundary.
@@
Graphene is typically manually exfoliated onto cleaved chips of crystalline material (such as silicon), each with an area of a few cm. The eXfoliator exfoliates over a 50x50 mm square (25 cm area, 5-20x that in a typical manual exfoliation) on a 76 mm diameter Si/SiO wafer. The increased scale was chosen in tandem with the purchase of a microscope with a motorized stage, capable of imaging contiguous areas as large as 8 cm on a side. To leverage this capability, a computer-vision algorithm was developed to reliably identify images containing monolayer graphene flakes by empirically relating local substrate color to monolayer appearance and identifying contiguous clusters of pixels, allowing for more efficient evaluation of exfoliations and reducing both labor and time required to locate the largest monolayers on the substrate. The imaging and flake identification will be described elsewhere, but similar techniques have been previously demonstrated[23].
In an ideal exfoliation of a van der Waals material, single-crystal flakes of the desired thickness – often, as here, monolayer – are cleaved from thicker crystals on the populated tape by adhesion to the substrate. The resulting size, integrity, cleanliness, and number density of these flakes on the substrate thus characterize the success of the exfoliation attempt.
@@ II Materials and Methods
The Scotch tape exfoliation process described here entails creating graphite-covered tape, applying the populated tape to an oxidized silicon substrate, annealing the tape-silicon system, and subsequently peeling the tape from the silicon. The silicon substrate must then be searched to locate resulting monolayers, which can be picked up and used to create graphene-based devices.
@@ II.1 The eXfoliator
The eXfoliator uses three perpendicular linear stepper motors (Fig. 1B, Parts 3,9) mounted on a metal frame (Parts 1-2,4-8) to move an extended arm (Parts 10-16) above a remotely controlled hot plate (Part 17) with a custom machined aluminum working surface (Parts 18-21). The machine itself occupies a nearly cubic volume of 46x54x51 cm. The motors, force gauge, and hot plate are connected to an external computer running a custom GUI to facilitate operation, including both manual and scripted control with recipes that can be standardized.
The hot plate used in this work was initially selected for its advertised heating and weighing functions with digital control; however, the weighing function was determined to be unreliable, necessitating incorporating a separate force gauge. Additionally, the time to reach temperature set points was substantially increased relative to the hot plate’s initial performance due to the additional thermal mass and surface area of the plate adapter and wafer pedestal (Parts 18-19). For future implementations of the eXfoliator design, we suggest considering other hot plates with increased heating power to more quickly reach temperature set points; a minor redesign of Part 18 will be required to accommodate a different working surface.
As shown in Figure 1B, the presented design was modified to improve cooling, thereby more closely approximating a common method of manual exfoliation where the tape-silicon sandwich is removed from the hot plate and allowed to cool within seconds prior to peeling. However, even the modified version took nearly 25 minutes to cool from 105C to below 30C; incorporating active cooling may prove beneficial.
@@ II.2 Graphene Exfoliation
In Scotch tape exfoliation, natural graphite crystals are repeatedly adhered to adhesive tape and cleaved apart to populate a large area of the tape with successively thinner graphite crystals; as this process roughly doubles the graphite-populated area of tape with each cleave, a 25 cm area may be fully populated in as few as 8 cleaves from a 9 mm starting crystal. The eXfoliator was designed to be able to populate two 25 cm areas of tape simultaneously, but the process proved too inefficient compared to manually populating the tape. Therefore, for all results reported here, we manually populated an initial ‘mother tape’, which we then used to populate a 25 cm ‘daughter tape’ to be used in the eXfoliator, a process analogous to that commonly used in manual exfoliation. The materials used by the eXfoliator in this paper are also commonly used in manual exfoliation: 1" Scotch magic tape, ”flaggy” graphite flakes purchased from NGS Trading & Consulting, and 76mm silicon wafers topped with 300nm wet silicon oxide purchased from NovaWafers. However, the eXfoliator’s design does not limit it to these materials. We envision that the eXfoliator can be easily used to exfoliate most vdW materials, including hBN and transition metal dichalcogenides, onto a wide variety of substrates using a wide variety of adhesive surfaces to hold the initial material.
During operation, a pristine 76 mm Si/SiO wafer is clamped in place on the wafer holder (Part 19) by Part 20. The daughter tape is mounted to the tape plate (Part 16) and populated in place; the tape plate is then mounted to the moving arm using Part 15, with four intervening wave springs to increase uniformity of contact as pressure is applied. The tape plate is positioned over the target substrate using the GUI, and the tape can then be applied to the substrate at a desired pressure, , annealed at a desired temperature, , and peeled at a desired rate, , at an angle in the Y-Z plane, defined as the angle formed by the tape with respect to the Y-axis as the tape is peeled from the substrate. Typical usage used a range of peel speeds and anneal temperatures, but targeted a 30 kPa application pressure, annealing over the time required to cool from the set temperature to within 5C of the desired peeling temperature of 25C, and a peel angle that was varied continuously from 60 to 80 over each exfoliation.
@@
| 60 | 170 | 60-80 | 30-170 |
After the tape is peeled, the wafer with exfoliated graphene is transferred to the motion stage of a Leica DM6 M microscope. The area on the wafer where exfoliation is performed is then scanned at 10x magnification with a flat-field correction to equalize brightness, and constant values for exposure, gain, white balance, gamma, color saturation, and illumination intensity. A focus map for the scan area is made using between 10 and 50 manually determined XYZ values as focal points across the wafer’s surface. Each scan generates approximately 3200 24-bit 20MP images per wafer to be analyzed by a flake-finding algorithm we designed for this purpose[25].

### v2

@@ Abstract
Over the past two decades, graphene has been intensively studied because of its remarkable mechanical, optical, and electronic properties. Initial studies were enabled by manual “Scotch Tape” exfoliation; nearly two decades later, this method is still widely used to obtain chemically-pristine flakes of graphene and other 2D van der Waals materials. Unfortunately, the yield of large, pristine flakes with uniform thickness is inconsistent. Thus, significant time and effort are required to exfoliate and locate flakes suitable for fabricating multilayer van der Waals heterostructures. Here, we describe a relatively affordable tabletop device (the “eXfoliator”) that can reproducibly control key parameters and largely automate the exfoliation process. In a typical exfoliation run, the eXfoliator produces 3 or more large (m2) high-quality graphene monolayer flakes, allowing new users to produce such flakes at a rate comparable to manual exfoliation by an experienced user. We use an automated mapping system and a computer vision algorithm to locate candidate flakes. Our results provide a starting point for future research efforts to more precisely identify which parameters matter for the success of exfoliation, and to optimize them.
@@ I Introduction
Monolayer graphene, a single layer of carbon atoms arranged in a honeycomb lattice with an atomic spacing of 1.42Å, has remarkable mechanical, optical, and electronic properties. Graphene has recently emerged as an exciting basis for studying strongly-correlated electron physics. Common stacking orders of multi-layer graphene, such as Bernal bilayer and rhombohedral trilayer and pentalayer can display superconductivity [1, 2, 3] and orbital magnetism[4, 5]. In addition to these configurations exfoliated from natural graphite, monolayer graphene can be artificially stacked with interlayer twists to form moiré materials with a large structural design space, including configurations with flat electronic bands and non-trivial band topology. Superconductivity and orbital magnetism[6, 7], and fractional Chern insulators[8, 9], have been discovered in various graphene-based moiré materials. Motivated by the rich set of target structures based on this single material, we focus our discussion and experimental demonstration on graphene, but we anticipate our methods and workflow will apply more broadly to the burgeoning family of exfoliatable van der Waals layered materials.
@@
In the Scotch tape exfoliation method introduced by Geim and Novoselov in 2004[10], natural or man-made graphite crystals –- which may consist of – graphene layers arranged in large crystallographic domains –- are repeatedly adhered to adhesive tape and cleaved apart to populate large areas of the tape with increasingly thin graphite crystals. The graphite-populated surface of the tape is then applied to a Si substrate coated with a thin SiO2 layer, before the tape is peeled away. Graphene flakes with a distribution of thicknesses ranging down to just one layer cleave from the graphite and are left on the substrate.
@@
Other techniques for producing large-area monolayers from bulk graphite, including electrostatic exfoliation, gold-tape exfoliation, intercalation, and sonication continue to be developed, and graphene can also be grown over large areas by chemical vapor deposition (CVD)[11, 12, 13, 14]. Each of these techniques has its own advantages, as discussed below, but Scotch tape exfoliation remains pre-eminent when seeking low disorder graphene-based structures in a research – as opposed to industrial production – environment.
@@
Electrostatic exfoliation uses high potentials between two metallic electrodes to directly cleave graphite into mono- and few-layer graphene flakes, but often results in smaller flakes with higher numbers of defects compared to flakes obtained by tape exfoliation, especially in monolayers[11].
@@
Metal-assisted or metal-tape exfoliation can produce large monolayers of transition metal dichalcogenides materials or graphene, starting from a bulk crystal or an epitaxial thin film[15, 16, 12]. These techniques depend on strongly adhering the target monolayer to a metal film. The monolayer is subsequently transferred to a new substrate and the metal film chemically etched away. Questions about whether metal and etchant can be fully removed without damaging the monolayer film have so far limited adoption of these approaches in research aimed at correlated electron states. Sonication and intercalation to extract monolayer graphene produce large numbers of flakes but leave adsorbates on the graphene and often produce flakes smaller than 1 micron[12, 17].
@@
Large-area growth techniques have produced promisingly large single-crystal graphene monolayers on a variety of substrates[18, 19, 13, 14]. Transfer of these monolayers to silicon has typically involved solvents and sometimes etchants. Dry transfer, most suitable for seeking exotic electronic states, has been rare for these grown films; despite notable successes in oxidizing a copper growth substrate then picking up graphene using exfoliated hBN[20], this method has not been widely adopted.
Some of these techniques are popular in particular contexts, but as noted above tape exfoliation remains dominant in experiments aimed at studying correlated electronic states. We set out to design a tabletop machine to reduce the time researchers must spend on tape exfoliation and to make the process’s yield more consistent. Our metrics for success include the production rate, number density, size, and disorder of flakes deposited.
@@
We hypothesized that if the key parameters of the exfoliation process could be identified and controlled, optimal values could be determined through experiments using a purpose-built tool, and then reproduced in successive exfoliations with that tool. When surveyed, researchers in our lab reported that manually exfoliating graphene on Si/SiO2 chips, including the initial tape population, takes between 1 and 6 minutes per square centimeter of exfoliated area. This time depends on an individual researcher’s chosen recipe, which can vary in the area prepared, technique for populating the tape, technique for pressing tape against substrate, anneal temperature and duration, and peel speed and angle. Given the anecdotally reported variation in both exfoliation technique and yield, we hypothesized that translating exfoliation to a semi-automated standardized process run on a machine could enable a researcher to produce more large single-crystal freestanding monolayers than a series of manual exfoliations in a comparable amount of time. The search for good process parameters need only be carried out once for each material, so any operator might be able to achieve consistent results after a shorter learning period than that required for manual exfoliation. A machine could also reach process parameters outside the range that a human researcher could readily and consistently implement manually (e.g., peeling at 0.1 mm/s or maintaining a chip at 80∘C during peeling), possibly enabling a machine to achieve better results than manual exfoliation in the hands of even an experienced practitioner. Finally, such a machine would free the operator to pursue other work while running an otherwise tedious or time-consuming exfoliation recipe.
@@
In fact, based on similar considerations, others have already implemented automated exfoliation systems. A prominent example of automated exfoliation is Brookhaven’s Quantum Material Press (QPress)[21], a sophisticated cluster of almost fully automated tools capable of exfoliating, cataloguing, stacking, and storing 2D materials in an inert atmosphere. QPress has successfully automated most of the tape exfoliation process, is accessible through user proposals, and may eventually provide flakes to many collaborators. In the opposite limit, a simple automated exfoliation tool can consist of a single linear motor integrated with a hot plate[22] to control peel speed and anneal temperature.
@@
The tool presented in this paper (the “eXfoliator”) was designed to fill a middle ground, as a user-friendly machine capable of controlling a range of likely-relevant parameters during exfoliation, balancing utility with simplicity of design and construction.
The parameters of Scotch tape exfoliation targeted by the design are peel angle, peel rate, application pressure, temperature of the substrate at all stages, and dwell time under pressure or at temperature. Other relevant parameters, such as preparation of the SiO2 surface by oxygen plasma cleaning, are known to have a significant effect but are not explicitly controlled by the eXfoliator[23].
@@
Estimates of graphite’s interlayer van der Waals binding energy vary, but most calculated values for AB, AAA, and ABC-stacked graphite range from 40-60 meV/atom[24]. For comparison, the thermal energy of graphite at 300 (400) K is approximately 1300 (2400) J/mol, corresponding to 13.5 (25) meV/atom[25]. Because the thermal energy at accessible temperatures – as limited by the melting point of the tape’s backing – does not approach the binding energy threshold, we assume tuning the annealing or peeling temperature primarily affects the viscoelastic behavior of the polymer tape adhesive.
@@
From the time-temperature superposition principle of polymer flow[26], we believe adjusting the annealing temperature, application pressure, or dwell time allows the viscoelastic polymer tape adhesive to flow more freely over the graphite flakes and allow better contact before being cooled and peeled. Likewise, we assume adjusting peel speed tunes the viscoelastic behavior of the tape adhesive instead of significantly affecting the fundamental interlayer interactions of the graphene layers. Adjusting the peel angle is believed to change both the radius of curvature of the tape and the ratio of shear to normal forces at the interlayer boundary.
@@
Graphene is typically manually exfoliated onto substrates with an area of one to a few cm2 cut from a Si/SiO2 wafer. The eXfoliator exfoliates over a 50x50 mm square (25 cm2 area, 5-20x that in a typical manual exfoliation) on a 76 mm diameter Si/SiO2 wafer. To take advantage of this increased scale we purchased a microscope with a motorized stage capable of imaging contiguous areas as large as 8 cm on a side, and we developed a computer-vision algorithm to reliably identify images containing monolayer graphene flakes. The algorithm empirically relates local substrate color to monolayer appearance, and identifies contiguous clusters of pixels, allowing efficient evaluation of exfoliation output and reducing both labor and time required to locate the largest monolayers on the substrate. The imaging and flake identification will be described elsewhere, but similar techniques have been previously demonstrated[27, 28].
@@ II Materials and Methods
The Scotch tape exfoliation process described here entails creating graphite-covered tape, applying the populated tape to an oxidized silicon substrate, annealing that substrate, and subsequently peeling the tape from the silicon. The silicon substrate must then be searched to locate resulting monolayers, which can be picked up and used to create graphene-based devices.
@@ II.1 The eXfoliator
A schematic of the eXfoliator is shown in Fig. 1. The figure caption points the reader to descriptions and CAD files for all parts. Three perpendicularly-oriented linear stepper motors (Fig. 1B, Parts 3,9) mounted on a metal frame (Parts 1-2,4-8) to move an extended arm (Parts 10-16) above a remotely controlled hot plate (Part 17) with a custom machined aluminum working surface (Parts 18-21). The machine itself occupies a nearly cubic volume of 46x54x51 cm. The motors, force gauge, and hot plate are connected to an external computer running a custom GUI that enables either manual control or scripting with recipes that can be standardized.
We selected the hot plate based on its computer control interface and its built-in weighing scale to measure the force with which the tape was being pressed onto the substrate; however, the weighing function proved unreliable, so we incorporated a separate force gauge. Adding a plate adapter and wafer pedestal (Parts 18-19) substantially increased the time needed to reach temperature set points. For future implementations of the eXfoliator design, we suggest considering other hot plates with increased heating power to more quickly reach temperature set points; a minor redesign of Part 18 would be required to accommodate a different working surface.
As shown in Figure 1B, we added fins to speed cooling, aiming to approximate a common method of manual exfoliation where the tape-silicon sandwich is removed from the hot plate and allowed to cool within seconds prior to peeling. However, even the modified version took nearly 25 minutes to cool from 105∘C to below 30∘C; incorporating active cooling may prove beneficial.
@@ II.2 Graphene Exfoliation
In Scotch tape exfoliation, natural or synthetic graphite crystals are repeatedly adhered to adhesive tape and cleaved apart to populate a large area of the tape with successively thinner graphite crystals; as this process roughly doubles the graphite-populated area of tape with each cleave, a 25 cm2 area may be fully populated in as few as 8 cleaves from a 9 mm2 starting crystal. The eXfoliator was designed to be able to populate two 25 cm2 areas of tape simultaneously, but the process proved inefficient compared to manually populating the tape. Therefore, for all results reported here, we manually populated an initial ‘mother tape’, which we then used to populate a 25 cm2 ‘daughter tape’ to be used in the eXfoliator, a process analogous to that commonly used in manual exfoliation. The materials used by the eXfoliator in this paper are also commonly used in manual exfoliation: 1" Scotch magic tape, “flaggy” graphite flakes purchased from NGS Trading & Consulting, and 76mm silicon wafers topped with 300nm wet silicon oxide purchased from NovaWafers. However, the eXfoliator’s design does not limit it to these materials. We envision that the eXfoliator can be easily used to exfoliate most vdW materials, including hBN and transition metal dichalcogenides, onto a wide variety of substrates using a wide variety of adhesive surfaces to hold the initial material.
During operation, a pristine 76 mm Si/SiO2 wafer is clamped in place on the wafer holder (Part 19) by Part 20. The tape plate (Part 16) is placed on a bench top with the tape-mounting surface face-up. Two side-by-side 1” strips of tape are mounted, sticky side up, to the tape plate. These serve as the daughter tape. The daughter tape is populated by gently pressing the mother tape to one half, before manually peeling it at a steep angle (typically, >120∘); the process is then repeated with the same mother tape on the other half. The tape plate is then mounted to the moving arm using Part 15, with four intervening parallel wave springs to increase uniformity of contact as pressure is applied. The tape plate is positioned over the target substrate using the GUI, and the tape can then be applied to the substrate at a desired pressure, , annealed at a desired temperature, , and peeled at a desired linear speed , at an angle in the Y-Z plane, defined as the angle formed by the tape with respect to the Y-axis as the tape is peeled from the substrate, assuming the tape remains unstretched. We have used a range of tape application temperatures and peel speeds, but targeted an initial application pressure of 30 kPa, releasing one side of the tape and removing the applied pressure as the wafer cooled from the set temperature to within 5∘C of the desired peeling temperature of 25∘C. The linear motion of the plate during peeling resulted in the peel angle varying continuously from 60∘ to 80∘ over each exfoliation, as calculated from the geometry of the system.
@@
| 60∘ | 170∘ | 60-80∘ | 30-170∘ |
After the tape is peeled, the wafer with exfoliated graphene is transferred to the motion stage of a Leica DM6 M microscope. The area on the wafer where exfoliation is performed is then imaged in an array of photos through a 10x objective. Exposure, gain, white balance, gamma, color saturation, and illumination intensity are kept constant. A flat-field correction equalizes brightness across each image. A focus map for the scan area is made using between 10 and 50 manually determined XYZ values as focal points across the wafer’s surface. Scanning each wafer generates approximately 3200 24-bit 20MP images to be analyzed by a flake-finding algorithm we designed for this purpose. Though we designed this algorithm independently, we later encountered a similar but more sophisticated method in the recent literature[28].